OpenStudy (kainui):

Find a function that is its own third derivative but not its own first or second derivative.

OpenStudy (anonymous):

would u plz state your question in mathematical words :) thank you

OpenStudy (kc_kennylau):

y'''=y Auxiliary equation: r^3-1=0, (r-1)(r^2+r+1)=0, r=1 or r^2+r+1=0. <sidenote>r^2+r+1=0, r=\(\dfrac{-1\pm\sqrt{-3}}2\)</sidenote>

OpenStudy (kc_kennylau):

Too complicated, I give up

OpenStudy (kc_kennylau):

P.S. It's now 23:53 in my timezone

OpenStudy (kainui):

No way, that's the fun of it, I think kenny pretty much got it, it's not really supposed to be insane difficulty lol.

OpenStudy (shamil98):

you learned the topic of second-order diff equations pretty quick kc , nice job

OpenStudy (kainui):

y=e^(e^(i*2pi/3)) I think it looks cooler like this.